# Quasiparticle-induced decoherence of a driven superconducting qubit

### v1

@@ Abstract
A recently developed strategy for counteracting the QP tunneling is known as gap engineering [26, 13, 27]. It amounts to implementing qubits with a difference $\delta\Delta=\Delta_{R}-\Delta_{L}$ in gap values across the JJ, see Fig. 1(a). The gap difference $\delta\Delta$ acts as a potential barrier for QPs. If $\delta\Delta/\hbar$ exceeds the qubit frequency $\omega_{\rm q}$—and QPs are “cold” [12, 13]—then QPs are unable to tunnel. This blocks the QP channel of the energy relaxation. The success of gap engineering is exemplified by a dramatic suppression of the effects of radiation impacts on qubit arrays [27].
@@
The term ^$\hat{\mathcal{H}}_{\rm T}(t)$ in Eq. (1) describes tunneling of QPs across the two junctions in the presence of a microwave drive  [^1] A similar model was recently considered in Ref. 44 in the context of QP effects in superconducting cat qubits.:
@@
where $D_{0}(\omega)=\omega^{2}_{\rm q}(0)/[\omega^{2}-\omega_{\rm q}^{2}(0)]$ is the qubit response function. The second term in the curly bracket originates from the parametric oscillation of the qubit frequency at $2\omega_{\rm d}$ [second diagram in Fig. 2(b)]. For $\omega_{\rm d}\sim\omega_{\rm q}$, it is of the same order as the first term. According to Eq. (11), the rate of the two-photon process is proportional to $S_{+}$; this needs to be contrasted with the rate of the one-photon process which is $\propto S_{-}\ll S_{+}$. In addition to featuring a larger structure factor, the two-photon process also remains effective in a wider range of drive frequencies, $\omega_{\rm d}>(\delta\Delta/\hbar-\omega_{\rm q})/2$. The rates of one- and two-photon processes have the same order of magnitude at $\Phi=0$ and $\omega_{\rm ac}$ relevant for the transmon readout [34]. In Ref. 34, we generalize Eq. (11) to arbitrary $\Phi$.
Excitation process — Not only can QPs relax a driven qubit, they can also lead to its excitation. To exemplify this phenomenon, we consider a $0\rightarrow 1$ qubit transition; we focus only on a one-photon process here deferring the discussion of a two-photon process to Ref. 34. A direct counterpart of Eq. (8) is
@@
Equation (14) generalizes the result of Ref. 35 from a single-photon case to the multi-photon one. The rate of qubit excitation, $\tilde{\Gamma}^{(n)}_{0\rightarrow 1}$, can be obtained by replacing $\omega_{\rm q}(\Phi)\rightarrow-\omega_{\rm q}(\Phi)$ in the argument of $\tilde{S}_{\pm p(n)}$ in Eq. (14).
Figure 1(c) shows the dependence of qubit excitation and relaxation rates on the drive frequency at $\Phi=0$. Its main feature is that $\tilde{\Gamma}_{0\rightarrow 1/1\rightarrow 0}$ change exponentially with $2\overline{\Delta}/(\hbar\omega_{\rm d})$ [36]. The thresholds which modulate the overall exponential trend stem from switches in the number of photons required to break a Cooper pair. The $n$-th threshold is located at $\omega_{\rm d}=(2\overline{\Delta}/\hbar-\omega_{if})/n$. The behavior near thresholds with even $n$ differs from that near thresholds with odd $n$. This parity effect stems from the different threshold behavior of $\tilde{S}_{+}$ and $\tilde{S}_{-}$, see Eqs. (14) and (16). In Ref. 34, we generalize Eq. (14) to arbitrary initial and final qubit states.
@@
The microwave-assisted QP tunneling can be suppressed at the hardware level. For example, the use of junctions with larger $\delta\Delta$ can set processes with small photon number below their thresholds. Another option is to implement QP traps that would keep the QPs away from the qubit’s JJs (see, e.g., Ref. 42).
Finally, let us comment on the pair-breaking process (second mechanism). This process represents a limitation on the performance of a high-frequency dispersive readout [40]. Ref. 40 showed that the readout fidelity can be improved by using a measurement tone with $\omega_{\rm d}\gg\omega_{\rm q}$. Our theory shows that $\omega_{\rm d}$ cannot be increased indefinitely; once the frequency becomes too high, the pair-breaking process starts to degrade the readout fidelity. Consider, for example, a $5~{}\rm GHz$ aluminum qubit that is read out at $\omega_{\rm d}/2\pi=60~{}\rm GHz$ [this is a scaled-up version of Ref. 40 setup]. The gap of aluminum is $\approx 50\,{\rm GHz}$; therefore, absorption of two photons can break a Cooper pair. Using Eq. (14) for $n=2$, we estimate that the fidelity is limited to $1-F\gtrsim 0.05$.
Note — Similar QP-induced decoherence mechanisms are also considered numerically in Ref. 43, which we became aware of during the preparation of the manuscript. Our results agree with each other.

### v2

@@ Abstract
A recently developed strategy for counteracting the QP tunneling is known as gap engineering [26, 13, 27, 28]. It amounts to implementing qubits with a difference $\delta\Delta=\Delta_{R}-\Delta_{L}$ in gap values across the JJ, see Fig. 1(a). The gap difference $\delta\Delta$ acts as a potential barrier for QPs. If $\delta\Delta/\hbar$ exceeds the qubit frequency $\omega_{\rm q}$—and QPs are “cold” [12, 13]—then QPs are unable to tunnel. This blocks the QP channel of the energy relaxation. The success of gap engineering is exemplified by a dramatic suppression of the effects of radiation impacts on qubit arrays [27].
@@
The term ^$\hat{\mathcal{H}}_{\rm T}(t)$ in Eq. (1) describes tunneling of QPs across the two junctions in the presence of a microwave drive  [^1] A similar model was recently considered in Ref. 49 in the context of QP effects in superconducting cat qubits.:
@@
where $D_{0}(\omega)=\omega^{2}_{\rm q}(0)/[\omega^{2}-\omega_{\rm q}^{2}(0)]$ is the qubit response function. The second term in the curly bracket originates from the parametric oscillation of the qubit frequency at $2\omega_{\rm d}$ [second diagram in Fig. 2(b)]. For $\omega_{\rm d}\sim\omega_{\rm q}$, it is of the same order as the first term. According to Eq. (11), the rate of the two-photon process is proportional to $S_{+}$; this needs to be contrasted with the rate of the one-photon process which is $\propto S_{-}\ll S_{+}$. In addition to featuring a larger structure factor, the two-photon process also remains effective in a wider range of drive frequencies, $\omega_{\rm d}>(\delta\Delta/\hbar-\omega_{\rm q})/2$. The rates of one- and two-photon processes have the same order of magnitude at $\Phi=0$ and $\omega_{\rm ac}$ relevant for the transmon readout [35]. In Ref. 35, we generalize Eq. (11) to arbitrary $\Phi$.
Excitation process — Not only can QPs relax a driven qubit, they can also lead to its excitation. To exemplify this phenomenon, we consider a $0\rightarrow 1$ qubit transition; we focus only on a one-photon process here deferring the discussion of a two-photon process to Ref. 35. A direct counterpart of Eq. (8) is
@@
Equation (14) generalizes the result of Ref. 36 from a single-photon case to the multi-photon one. The rate of qubit excitation, $\tilde{\Gamma}^{(n)}_{0\rightarrow 1}$, can be obtained by replacing $\omega_{\rm q}(\Phi)\rightarrow-\omega_{\rm q}(\Phi)$ in the argument of $\tilde{S}_{\pm p(n)}$ in Eq. (14).
Figure 1(c) shows the dependence of qubit excitation and relaxation rates on the drive frequency at $\Phi=0$. Its main feature is that $\tilde{\Gamma}_{0\rightarrow 1/1\rightarrow 0}$ change exponentially with $2\overline{\Delta}/(\hbar\omega_{\rm d})$ [37]. The thresholds which modulate the overall exponential trend stem from switches in the number of photons required to break a Cooper pair. The $n$-th threshold is located at $\omega_{\rm d}=(2\overline{\Delta}/\hbar-\omega_{if})/n$. The behavior near thresholds with even $n$ differs from that near thresholds with odd $n$. This parity effect stems from the different threshold behavior of $\tilde{S}_{+}$ and $\tilde{S}_{-}$, see Eqs. (14) and (16). In Ref. 35, we generalize Eq. (14) to arbitrary initial and final qubit states.
@@
The microwave-assisted QP tunneling can be suppressed at the hardware level. For example, the use of junctions with larger $\delta\Delta$ can set processes with small photon number below their thresholds. Another option is to implement QP traps that would keep the QPs away from the qubit’s JJs (see, e.g., Ref. 43).
Finally, let us comment on the pair-breaking process (second mechanism). This process represents a limitation on the performance of a high-frequency dispersive readout [41]. Ref. 41 showed that the readout fidelity can be improved by using a measurement tone with $\omega_{\rm d}\gg\omega_{\rm q}$. Our theory shows that $\omega_{\rm d}$ cannot be increased indefinitely; once the frequency becomes too high, the pair-breaking process starts to degrade the readout fidelity. Consider, for example, a $5~{}\rm GHz$ aluminum qubit that is read out at $\omega_{\rm d}/2\pi=60~{}\rm GHz$ [this is a scaled-up version of Ref. 41 setup]. The gap of aluminum is $\approx 50\,{\rm GHz}$; therefore, absorption of two photons can break a Cooper pair. Using Eq. (14) for $n=2$, we estimate that the fidelity is limited to $1-F\gtrsim 0.05$.
The mechanisms we uncover are relevant for applications of microwave drives to superconducting circuits beyond transmon qubits [44, 45, 46, 47].
Note — Similar QP-induced decoherence mechanisms are also considered numerically in Ref. 48, which we became aware of during the preparation of the manuscript. Our results agree with each other.
@@
The research of P.D.K. and T.C. was sponsored by the Army Research Office (ARO) under grants no.W911NF-22-1-0053 and W911NF-23-1-0051, by DARPA under grant no. HR0011-24-2-0346, by the U.S. Department of Energy (DoE), Office of Science, National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage (C2QA) under contract number DE-SC0012704. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the ARO, DARPA, DoE, or the US Government. The US Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.